OpenStudy (anonymous):

I need hep with an algebra 2 question: given that f(x)=3x^2-4 and g(x)=2x-6, what is g(f(2))?

OpenStudy (anonymous):

\[\Large f(x)=3x^2-4\] \[\Large f(2)=3\cdot 2^2-4\] \[\Large f(2)=12-4\] \[\Large f(2)=8\] \[\Large g(x)=2x-6\] \[\Large g(f(2))=2(f(2))-6\] \[\Large g(f(2))=2\cdot 2-6\] \[\Large g(f(2))=4-6\] \[\Large g(f(2))=-2\]

OpenStudy (kinggeorge):

Kreshnik's solution is good up until he plugs in for \(g(f(2))\) Since \(f(2)=8\)\[\Large g(f(2))=g(8)\]\[\Large g(f(2))=2\cdot8-6\]\[\Large g(f(2))=10\]

OpenStudy (anonymous):

auff... LOL @KingGeorge Thanks for correcting me.. I wrote there: \[\LARGE f(2)=8\] but when I plugged in I changed that :( .